# Towards a Surgeon-in-the-Loop Ophthalmic Robotic Apprentice using Reinforcement and Imitation Learning

### v2

@@ Abstract
Robot-assisted surgical systems have demonstrated significant potential in enhancing surgical precision and minimizing human errors. However, existing systems cannot accommodate individual surgeons’ unique preferences and requirements. Additionally, they primarily focus on general surgeries (e.g., laparoscopy) and are unsuitable for highly precise microsurgeries, such as ophthalmic procedures. Thus, we propose an image-guided approach for surgeon-centered autonomous agents that can adapt to the individual surgeon’s skill level and preferred surgical techniques during ophthalmic cataract surgery. Our approach trains reinforcement and imitation learning agents simultaneously using curriculum learning approaches guided by image data to perform all tasks of the incision phase of cataract surgery. By integrating the surgeon’s actions and preferences into the training process, our approach enables the robot to implicitly learn and adapt to the individual surgeon’s unique techniques through surgeon-in-the-loop demonstrations. This results in a more intuitive and personalized surgical experience for the surgeon while ensuring consistent performance for the autonomous robotic apprentice. We define and evaluate the effectiveness of our approach in a simulated environment using our proposed metrics and highlight the trade-off between a generic agent and a surgeon-centered adapted agent. Finally, our approach has the potential to extend to other ophthalmic and microsurgical procedures, opening the door to a new generation of surgeon-in-the-loop autonomous surgical robots. We provide an open-source simulation framework for future development and reproducibility at time of publication.
@@ II Background and Related Work
II-A Automation of Surgical Robots
Surgery, as a medical specialty, uses operative manual and instrumental techniques applied to the human body to treat a patient’s injury or pathological condition. Such microscopic procedures are performed manually by experienced surgeons with the help of a surgical microscope and require, in addition to trained hand-eye coordination and the planned procedure, machines and technology to make time-critical and risky interventions feasible. More recently, with the introduction of surgical robotic systems, the operating surgeon manually configures surgical robotic imaging microscopes to ensure optimal support through the imaging properties of the visual device directed at the surgical zone. However, imaging requirements in terms of positioning, viewing, and focus plane may change throughout the operation, making it necessary for surgeons to adapt the configuration of the surgical microscope. Since adjustments must be performed manually, a short-term interruption of the operation becomes necessary. Individual risk considerations are made in each case as to whether visual section correction is effective in the present surgical situation, which could negatively affect the outcome of the surgery if taken incorrectly. Hence, there have been recent advances in the hand-free configuration of the microscope, such as mouth-controlled joysticks or foot control with the help of pedals [16, 17, 18]. However, these approaches are impractical and allow too few degrees of freedom to meet the complex requirements for the various settings of the microscopic frame, unlike a fully automated approach as presented in this work.
@@ II-C Imitation Learning
One of the methodologies used to solve the dilemma of defining the reward function with RL is imitation learning (IL) or learning from demonstrations, focusing on training agents to mimic expert behavior. Unlike RL, where agents learn through trial and error, IL uses expert demonstrations as guidance. In this approach, an agent learns a policy by observing and mimicking the actions performed by human or expert demonstrators. By learning from preexisting knowledge and demonstrations, imitation learning offers several advantages, such as faster convergence and improved sample efficiency compared to pure reinforcement learning [25, 26]. It finds applications in various domains, including robotics [27, 28, 29, 30], autonomous driving [31, 32, 33], and virtual simulation [34]. Imitation learning algorithms, such as Behavioral Cloning [35], Inverse Reinforcement Learning (IRL) [36], and Generative Adversarial Imitation Learning (GAIL) [37], provide frameworks for effectively transferring expert knowledge to autonomous agents. However, imitation learning also faces challenges such as handling distribution shifts between expert demonstrations and agent exploration, generalization to novel situations, and addressing the trade-off between imitation accuracy and exploration for robust decision making [25]. Therefore, our work hierarchically combines imitation and reinforcement learning to achieve a trade-off between the two methodologies within adequate performance levels.
@@ III-A Simulation Environment
The eye model used in this study comprises a comprehensive representation of its various anatomical components, including the cornea, iris, veins, arteries, and optic nerves. Each component is modeled to capture its unique characteristics and structural intricacies. The cornea, which acts as the transparent outermost layer, is reconstructed to emulate its curved shape and refractive properties. The iris, responsible for regulating the amount of light entering the eye, is depicted with its intricate patterns and radial muscle fibers. Voxelization techniques were employed to discretize the eye model into a three-dimensional grid, with each voxel measuring 0.01mm in size. This high resolution voxelization enables precise simulations of surgical procedures, such as cutting or incisions, and ensures accurate modeling of tissue interactions. The veins, arteries, and optic nerves, crucial for the eye’s vascular supply and visual information transmission, are represented to mimic their branching networks and connectivity within the eye. Therefore, our eye model provides a realistic and fine-grained simulation environment, enabling the agent to simulate and learn from intricate surgical maneuvers within the eye with a high degree of fidelity.

### v3

@@ Abstract
Robot-assisted surgical systems have demonstrated significant potential in enhancing surgical precision and minimizing human errors. However, existing systems cannot accommodate individual surgeons’ unique preferences and requirements. Additionally, they primarily focus on general surgeries (e.g., laparoscopy) and are unsuitable for highly precise microsurgeries, such as ophthalmic procedures. Thus, we propose an image-guided approach for surgeon-centered autonomous agents that can adapt to the individual surgeon’s skill level and preferred surgical techniques during ophthalmic cataract surgery. Our approach trains reinforcement and imitation learning agents simultaneously using curriculum learning approaches guided by image data to perform all tasks of the incision phase of cataract surgery. By integrating the surgeon’s actions and preferences into the training process, our approach enables the robot to implicitly learn and adapt to the individual surgeon’s unique techniques through surgeon-in-the-loop demonstrations. This results in a more intuitive and personalized surgical experience for the surgeon while ensuring consistent performance for the autonomous robotic apprentice. We define and evaluate the effectiveness of our approach in a simulated environment using our proposed metrics and highlight the trade-off between a generic agent and a surgeon-centered adapted agent. Finally, our approach has the potential to extend to other ophthalmic and microsurgical procedures, opening the door to a new generation of surgeon-in-the-loop autonomous surgical robots. We provide an open-source simulation framework for future development and reproducibility at https://github.com/amrgomaaelhady/CataractAdaptSurgRobot.
@@ II Background and Related Work
II-A Surgical Microscopy and Robotics
Surgery, as a medical specialty, uses operative manual and instrumental techniques applied to the human body to treat a patient’s injury or pathological condition. Such microscopic procedures are performed manually by experienced surgeons with the help of a surgical microscope and require, in addition to trained hand-eye coordination and the planned procedure, machines and technology to make time-critical and risky interventions feasible. More recently, with the introduction of surgical robotic systems, the operating surgeon manually configures surgical robotic imaging microscopes to ensure optimal support through the imaging properties of the visual device directed at the surgical zone. However, imaging requirements in terms of positioning, viewing, and focus plane may change throughout the operation, making it necessary for surgeons to adapt the configuration of the surgical microscope. Since adjustments must be performed manually, a short-term interruption of the operation becomes necessary. Individual risk considerations are made in each case as to whether visual section correction is effective in the present surgical situation, which could negatively affect the outcome of the surgery if taken incorrectly. Hence, there have been recent advances in the hand-free configuration of the microscope, such as mouth-controlled joysticks or foot control with the help of pedals [16, 17, 18]. However, these approaches are impractical and allow too few degrees of freedom to meet the complex requirements for the various settings of the microscopic frame [19], unlike a fully automated approach as presented in this work.
@@ II-C Imitation Learning
One of the methodologies used to solve the dilemma of defining the reward function with RL is imitation learning (IL) or learning from demonstrations, focusing on training agents to mimic expert behavior. Unlike RL, where agents learn through trial and error, IL uses expert demonstrations as guidance. In this approach, an agent learns a policy by observing and mimicking the actions performed by human or expert demonstrators. By learning from preexisting knowledge and demonstrations, imitation learning offers several advantages, such as faster convergence and improved sample efficiency compared to pure reinforcement learning [26, 27]. It finds applications in various domains, including robotics [28, 29, 30, 31, 32], autonomous driving [33, 34, 35], and virtual simulation [36]. Imitation learning algorithms, such as Behavioral Cloning [37], Inverse Reinforcement Learning (IRL) [38], and Generative Adversarial Imitation Learning (GAIL) [39], provide frameworks for effectively transferring expert knowledge to autonomous agents. However, imitation learning also faces challenges such as handling distribution shifts between expert demonstrations and agent exploration, generalization to novel situations, and addressing the trade-off between imitation accuracy and exploration for robust decision making [26]. Therefore, our work hierarchically combines imitation and reinforcement learning to achieve a trade-off between the two methodologies within adequate performance levels.
@@ III-A Simulation Environment
The eye model used in this study comprises a comprehensive representation of its various anatomical components, including the cornea, iris, veins, arteries, and optic nerves. Each component is modeled to capture its unique characteristics and structural intricacies. The cornea, which acts as the transparent outermost layer, is reconstructed to emulate its curved shape and refractive properties. The iris, responsible for regulating the amount of light entering the eye, is depicted with its intricate patterns and radial muscle fibers. Voxelization techniques were employed to discretize the eye model into a three-dimensional grid, with each voxel measuring 0.01mm in size. This high resolution voxelization enables precise simulations of surgical procedures, such as cutting or incisions, and ensures accurate modeling of tissue interactions. The veins, arteries, and optic nerves, crucial for the eye’s vascular supply and visual information transmission, are represented to mimic their branching networks and connectivity within the eye. Moreover, deformation in the eye model through cutting and incision is simulated by removing the affected voxels to show the layers behind it (e.g., blood vessels) which mimics tissue displacement in real-world physics as seen in [44]. Therefore, our eye model provides a realistic and fine-grained simulation environment, enabling the agent to simulate and learn from intricate surgical maneuvers within the eye with a high degree of fidelity.